# Geometric theory of quark confinement

### v5

@@ Abstract
We present the nonperturbative solution of the loop equation in quenched QCD (one quark loop in full gluon vacuum, including nonplanar graphs). This solution relies on a specific local minimum of the Plateau problem—one that is additive over the closed parts of the bounding loop formed at self-intersections. This surface applies to large loops, leading to quark confinement via a factor $\exp\left(-\kappa S[C]\right)$ multiplying the perturbative Wilson loop $W_{pert}[C]$. Crucially, the confinement mechanism relies on the self-duality of the area derivative—a property that exists exclusively in four dimensions. This geometric constraint ensures stability only in four dimensions, distinguishing the resulting spectrum from standard string models, which are stable only in higher embedding dimensions. We compute the high-energy meson spectrum resulting from this novel confinement mechanism. The result is a usual linear Regge trajectory with the slope $\alpha^{\prime}=\frac{1}{2\pi\sigma}$ in our normalization of string tension $\sigma=2\sqrt{2}\kappa$. However, there are no string modes to be added to the spectrum in our solution, which amounts to the static linear potential for the quark pair. As we argue, the fluctuations of "flux tube" between quarks are accounted for in the gluon diagrams in $W_{pert}[C]$, and do not contribute to the confining force. This eliminates the problems of quantization of string in four dimensions.
@@ 1 Introduction
In this paper we consider the $N_{f}\to 0$ (quenched) valence limit of QCD at fixed $N_{c}$. The vacuum is pure Yang–Mills; quark currents are represented by a single closed valence quark loop whose propagation amplitude is multiplied by the Wilson loop $W[C]$ in the gluonic vacuum. In this limit, hadronic states cannot decay by creating additional $q\bar{q}$ pairs, and the meson spectrum is purely confining (rising to infinity). We approximate $W[C]$ by the Hodge–dual factor $\exp\left(-\kappa S[C]\right)$, neglecting subleading gluon–graph corrections in the large–loop limit. This factor, as we argue in this work, represents an exact solution of the MM chain of multi-loop equations [MMEq79, MM1981NPB, Mig83].
We specifically analyze a helical loop configuration (the Helicoid), which allows for an analytic WKB computation of the meson mass at large angular momentum, where the radius of the helix is large.
@@ 2 Self-dual area derivative and loop equation
Let us remind why we are looking for the self-dual minimal surface, as suggested in [migdal2025SQYMflow]. The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
@@
The loop equation for the Yang–Mills gradient flow was shown in [migdal2025SQYMflow] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
@@
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [migdal2025SQYMflow] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents the SD or ASD tensor. This choice would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
@@
It was heuristically claimed in the original MM papers [MM1981NPB, Mig83], and rigorously proven in the loop calculus [migdal2025SQYMflow], section 3, eqs (39), (40). Due to the Leibnitz property,
@@
We are looking for such a self-dual minimal surface which is additive at the self-intersecting loops. This condition will be discussed in detail below. It is cricial for the whole theory– due to this additivity the confining factor $\exp\left(-\kappa S[C]\right)$ is compatible with the whole chain of the loop equations.
@@
The full chain of loop equations for the Wilson loops with a finite number of colors $N_{c}$ [MM1981NPB, Mig83] has the same loop operator (2.3) as the Yang–Mills gradient flow, except there is no time derivative. In this paper, we are only considering the full chain of the multiloop equations for QCD, and not the Yang–Mills gradient flow equations. We use, however, the loop calculus developed in that paper, as it applies to the loop diffusion operator (2.3), which is the same in the multiloop MM chain and in the Yang–Mills gradient flow equation. It has the structure
@@
So, the kernel is parameterization-invariant. Therefore, the canonical standardization to $0,2\pi]$ used above is legitimate and additivity of $S[C]$ guaranties that the multiplicative dressing $W\mapsto W\,\exp\left(-\kappa S\right)$ preserves the MM factorization $\mathcal{L}W[C_{1}\!\circ\!C_{2}]=W[C_{1}]\otimes W[C_{2}]$ whenever $\mathcal{L}S=0$ (see (2.3) and [Mig83, migdal2025SQYMflow]).
@@ 2.2 Confining factor in the multiloop hierarchy
Let $\{W_{n}\}_{n\geq 1}$ denote the $n$-loop Wilson loop averages in pure $\mathrm{SU}(N_{c})$ Yang–Mills theory, satisfying the MM multiloop hierarchy [Mig83]. Let $S[C]$ be the geometric functional constructed above, with the properties
@@
The same loop operator $L$ in (2.3) appears on the left-hand side of the general QCD loop equations [Mig83] for arbitrary $N_{c}$. The difference between the planar and finite-$N_{c}$ cases lies only in the right-hand side: at a self-intersection, the planar factorization is replaced by a combination of one- and two-loop averages, but the kernel of the bilocal integral is the same.
@@
This equation is one link in the chain relating $W_{n}$ to $W_{n\pm 1}$; for finite $N_{c}$ the chain terminates due to the Fierz identity for $\mathrm{SU}(N_{c})$ [Mig83], a fact recently exploited in lattice studies of the loop equation [kazakovZheng2024].
@@
The parameter $\kappa$ is not determined by the chain of MM equations, nor by the initial value at the vanishing loop $W[0]=1$ (as $S[0]=0$). The boundary condition $W[\infty]=0$ only requires $\kappa\geq 0$. Therefore, this parameter is a free nonperturbative parameter, similar to the vacuum expectation value $\left\langle\operatorname{tr}F_{\mu\nu}^{2}\right\rangle$. The MM equations reflect the equations of motion of QCD, but not the property of the vacuum, so this ambiguity in the solutions of the equations of motion is inevitable: one must add some information regarding the properties of the gluon vacuum state. Later in this paper, we relate this parameter $\kappa$ to the slope of the Regge trajectories, which tells us that it is positive in the physical vacuum and has the same dimension as $\Lambda_{QCD}\sim M^{2}$. Ideally, in complete theory, one would have to relate it directly to $\Lambda_{QCD}$ times a universal dimensionless factor. This important task exceeds the scope of the present paper.
@@ 3 Hodge -dual minimal surface
Now, we are going to define the solvable functional $S[C]$ satisfying both the duality and additivity. We are following the general construction from [migdal2025SQYMflow], but here we correct an error made in that paper: we only need three components of the quaternionic field $X_{\mu}$, not four111the fourth component was worse than unnecessary, it canceled the correct area derivative, produced by the first three components. In the general proofs in that paper, this cancellation was overlooked. The intended results are restored by dropping the fourth component..
@@ 3.1 The surface
$\delta S=-2\int d\theta\frac{\delta C_{\mu}\Sigma_{\mu\nu}\partial_{\theta}C_{\nu}}{\sqrt{\Sigma^{2}}}$ (3.7)
Therefore, the Hodge duality of the area element $\Sigma$ with these boundary conditions for generic field $X^{a}_{\mu}(z,\bar{z})$ leads to the same Hodge duality of the area derivative.
This self-duality of area derivative makes our Hodge-dual minimal area a zero mode of the loop diffusion operator, leading to the solution of the MM loop equations [MM1981NPB, Mig83], as we argued in the next section.
@@
In the previous paper [migdal2025SQYMflow], we found exact solutions for the special case of a planar bounding loop and proved two inequalities. Unfortunately, four components of $X$ field instead of three were erroneously used in that work. In this paper, we find a complete solution for the Hodge-dual surface, superseding the results of previous work and correcting these errors.
@@ 3.3 The Virasoro Constraint and Uniformization
The existence of such a coordinate system for any surface topology is guaranteed by the Uniformization Theorem [poincare1907uniformisation]. This theorem ensures that we are always free to set $(f^{\prime})^{2}=0$ by a suitable diffeomorphism. In this gauge, the non-linear geometric area density simplifies exactly to the quadratic Lagrangian:
@@
$\frac{\delta S_{\chi}[C]}{\delta\sigma_{\mu\nu}}=2\frac{\Sigma_{\mu\nu}}{\sqrt{\Sigma_{\mu\nu}^{2}}};$ (3.26)
$\Sigma_{\mu\nu}=2(F_{\mu\nu}+\chi\ast F_{\mu\nu});$ (3.27)
@@
In presence of the Virasoro constraint (3.21), the problem becomes a challenging one, and it was investigated over the last centuries by great mathematicians, starting with Riemann and Hilbert. We give a brief summary of modern state of this theory in  A.
@@ 3.6 Symmetrization of the area by parity
One may then ask: why do we need the Hodge-dual surface at all? If the area derivative of the symmetrized Hodge-dual surface is equivalent (up to normalization) to that of the Nambu-Goto string, why not claim the Nambu-Goto string itself solves the loop equation as a zero mode? Direct calculation in [migdal2025SQYMflow] shows this is not possible. The area derivative of the conventional minimal area is well-defined, but it does not solve the loop equation: it produces in this equation a singular factor $\frac{\delta}{\delta^{2}+\epsilon^{2}}$, depending on the ratio of the two infinitesimal cutoff parameters involved in the definition of the loop variations. Depending on this ratio, we obtain an arbitrary number between zero and infinity instead of the required zero in the loop equation.
@@
$\frac{\delta S_{\chi}}{\delta\sigma}$ $\propto(F+\chi\ast F)/\sqrt{F^{2}};$
$F_{\mu\nu}$ $=\Im f_{\mu}\bar{f}_{\nu}$
@@ 3.10 Resulting Geometry
This solution represents (up to rotation with $\eta^{a}$ matrices) a general conformal mapping from the parameter domain $\xi$ directly to the physical plane $(X_{1},X_{2})$, with trivial coordinates in the transverse dimensions $X_{3},X_{4}$. The function $w(\xi)$ completely determines the shape of the planar loop. This planar solution matches the one found in [migdal2025SQYMflow] for planar loop.
@@
For this planar solution, we can explicitly verify the zero-mode condition. The area derivative for the planar ansatz reduces to the simple algebraic form $\delta S_{\chi}/\delta\sigma_{\mu\nu}\propto(F_{\mu\nu}+\chi\ast F_{\mu\nu})/\sqrt{F^{2}}$ which is manifestly self-dual. Therefore, the condition $\frac{\delta S_{\chi}}{\delta\sigma}=\chi\frac{\delta S_{\chi}}{\delta\sigma}$ holds pointwise. By the Bianchi identity (2.5), this implies $\mathcal{L}S[C]\equiv 0$ without reliance on the general loop calculus of Ref [4]. This confirms that the explicit planar solution is indeed an exact zero-mode of the loop operator. This is, however, a rather trivial zero mode, as the normalized tensor $(F_{\mu\nu}+\chi\ast F_{\mu\nu})/\sqrt{F^{2}}\propto\eta^{3,\chi}_{\mu\nu}$ is a universal constant tensor, with vanishing functional derivative.
@@ 4.1 Stationarity and Boundary Variation
The existence of the additive minimal solution is guaranteed by the results of [Douglas1931] and [Rado1930], who established that the disconnected solution minimizes the functional for separated or touching boundaries. Since the surface is an extremum in the full functional space, the bulk variation vanishes identically by the principle of stationary action.
@@
While a connected minimal surface (global minimum) may exist, the QCD Loop Equation requires the additive Goldschmidt solution to satisfy the factorization condition $S[C_{1}\cdot C_{2}]=S[C_{1}]+S[C_{2}]$. The topological barrier proven by [Gulliver1973] ensures that this additive solution is a local minimum, which is all we need for the proof of Hodge-duality.
@@
The continuity of the transition between a single loop with an infinitesimal neck and two touching loops (the self-intersecting limit) is mathematically grounded in the Bridge Principle for stable minimal surfaces [White1996]. This principle establishes that given two stable minimal surfaces bounded by curves $C_{1}$ and $C_{2}$, one can construct a connected minimal surface bounded by a single loop formed by joining $C_{1}$ and $C_{2}$ with a thin “bridge” of width $\epsilon$. As $\epsilon\to 0$, this connected surface converges smoothly (in the varifold sense) to the union of the original surfaces. Consequently, the Goldschmidt solution we utilize—which effectively treats the area as additive, $\mathcal{A}(C_{1}\cup C_{2})=\mathcal{A}(C_{1})+\mathcal{A}(C_{2})$—represents the correct continuous limit of the loop equation dynamics, avoiding the discontinuous jumps associated with the global minimization of the Plateau problem (such as the catenoid collapse). This ensures that the loop derivative remains well-defined across topology changes. Specifically, because the convergence is smooth in the varifold sense, the functional derivatives required for the loop equation (which probe the surface near the loop $C$) are continuous limits of the derivatives of the connected surface, justifying the application of the differential loop calculus to the additive Goldschmidt functional.
We assume (or conjecture, supported by [Douglas1931, Rado1930, White1996]) that the Goldschmidt/local-minimum branch defines a continuous additive functional on immersed loops with transverse self-intersections, and that the loop-space area derivative exists almost everywhere along each branch. Under this assumption, Theorem 2.1 applies.

### v6

# Geometric QCD I: The Hodge-Dual Surface and Quark Confinement

@@ Abstract
This is the first of two papers presenting a geometric framework for Planar QCD ($N_{c}\to\infty$). In this part, we establish the kinematic foundation of the theory by constructing the unique stable vacuum of the loop equation. We demonstrate that the Makeenko-Migdal loop equation admits a solution of the form $W[C]=W_{pert}[C]\exp\left(-\kappa S[C]\right)$, provided $S[C]$ is a specific minimal surface possessing a self-dual area derivative. We prove that such a surface exists and corresponds to the Hodge-dual projection of a minimal surface in $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$. Crucially, this confinement mechanism relies on the self-duality of the area derivative—a property that exists exclusively in four dimensions. This geometric constraint ensures stability only in $D=4$, distinguishing the resulting theory from standard string models which require higher critical dimensions. We relate the string tension parameter $\kappa$ to the gluon condensate via the Operator Product Expansion. The dynamical quantization of the Fermi string on this rigid surface and the resulting meson spectrum are derived in the companion paper [16].
@@ 1 Introduction
In this series of two papers, we propose a constructive approach to the long-standing problem of the string dual of Planar QCD. The solution requires addressing two fundamental distinct problems: the kinematics of the loop space (finding a stable vacuum geometry that satisfies the loop equations) and the dynamics of the flux tube (quantizing the excitations on this background to find the spectrum).
In this first paper (Part I), we solve the geometric problem. We consider the $N_{f}\to 0$ (quenched) valence limit of QCD at fixed $N_{c}$. The vacuum is pure Yang–Mills; quark currents are represented by a single closed valence quark loop whose propagation amplitude is multiplied by the Wilson loop $W[C]$ in the gluonic vacuum. We approximate $W[C]$ by the Hodge–dual factor $\exp\left(-\kappa S[C]\right)$, neglecting subleading gluon–graph corrections in the large–loop limit. This factor, as we argue in this work, represents a zero-mode solution of the MM chain of multi-loop equations [11, 12, 14].
The central result of this paper is the identification of the Hodge-dual minimal surface. We prove that this surface is the unique additive solution to the Plateau problem that possesses a self-dual area derivative, a property required to satisfy the Bianchi identity in loop space. This rigidity eliminates the Liouville instability inherent to random surface sums.
In the companion paper [16], we populate this rigid surface with internal Majorana fermions ("Elves"). We show that the resulting "Fermi String" theory reproduces the planar graphs of QCD and, crucially, leads to a finite effective action in Momentum Loop Space. The spectral problem is then solved in [16], yielding linear Regge trajectories.
@@ 2 Self-dual area derivative and loop equation
Let us remind why we are looking for the self-dual minimal surface, as suggested in [15]. The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
@@
The loop equation for the Yang–Mills gradient flow was shown in [15] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
@@
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [15] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents the SD or ASD tensor. This choice would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
@@
It was heuristically claimed in the original MM papers [12, 14], and rigorously proven in the loop calculus [15], section 3, eqs (39), (40). Due to the Leibnitz property,
@@
We are looking for such a self-dual minimal surface which is additive at the self-intersecting loops. This condition will be discussed in detail below. It is critial for the whole theory– due to this additivity the confining factor $\exp\left(-\kappa S[C]\right)$ is compatible with the whole chain of the loop equations.
@@
The full chain of loop equations for the Wilson loops with a finite number of colors $N_{c}$ [12, 14] has the same loop operator (2.3) as the Yang–Mills gradient flow, except there is no time derivative. In this paper, we are only considering the full chain of the multiloop equations for QCD, and not the Yang–Mills gradient flow equations. We use, however, the loop calculus developed in that paper, as it applies to the loop diffusion operator (2.3), which is the same in the multiloop MM chain and in the Yang–Mills gradient flow equation. It has the structure
@@
So, the kernel is parameterization-invariant. Therefore, the canonical standardization to $0,2\pi]$ used above is legitimate and additivity of $S[C]$ guaranties that the multiplicative dressing $W\mapsto W\,\exp\left(-\kappa S\right)$ preserves the MM factorization $\mathcal{L}W[C_{1}\!\circ\!C_{2}]=W[C_{1}]\otimes W[C_{2}]$ whenever $\mathcal{L}S=0$ (see (2.3) and [14, 15]).
@@
While the mathematical solution is exact, its identification with the QCD vacuum is a physical hypothesis supported by three facts:
  1. 1.
    it is related to the gluon condensate
  2. 2.
    it yields the correct Area Law (confinement)
  3. 3.
    it produces the correct linear Regge trajectories (particle spectrum)
@@ 2.2 Confining factor in the multiloop hierarchy
Let $\{W_{n}\}_{n\geq 1}$ denote the $n$-loop Wilson loop averages in pure $\mathrm{SU}(N_{c})$ Yang–Mills theory, satisfying the MM multiloop hierarchy [14]. Let $S[C]$ be the geometric functional constructed above, with the properties
@@
The same loop operator $L$ in (2.3) appears on the left-hand side of the general QCD loop equations [14] for arbitrary $N_{c}$. The difference between the planar and finite-$N_{c}$ cases lies only in the right-hand side: at a self-intersection, the planar factorization is replaced by a combination of one- and two-loop averages, but the kernel of the bilocal integral is the same.
@@
This equation is one link in the chain relating $W_{n}$ to $W_{n\pm 1}$; for finite $N_{c}$ the chain terminates due to the Fierz identity for $\mathrm{SU}(N_{c})$ [14], a fact recently exploited in lattice studies of the loop equation [9].
@@
The parameter $\kappa$ is not determined by the chain of MM equations, nor by the initial value at the vanishing loop $W[0]=1$ (as $S[0]=0$). The boundary condition $W[\infty]=0$ only requires $\kappa\geq 0$. Therefore, this parameter is a free nonperturbative parameter, similar to the vacuum expectation value $\left\langle\operatorname{tr}F_{\mu\nu}^{2}\right\rangle$. The MM equations reflect the equations of motion of QCD, but not the property of the vacuum, so this ambiguity in the solutions of the equations of motion is inevitable: one must add some information regarding the properties of the gluon vacuum state. Later in this paper, we relate this parameter $\kappa$ to the slope of the Regge trajectories, which tells us that it is positive in the physical vacuum and has the same dimension as $\Lambda^{2}_{QCD}\sim M^{2}$. Ideally, in complete theory, one would have to relate it directly to $\Lambda_{QCD}$ times a universal dimensionless factor. This important task will be addressed partly later in this paper, and elaborated deeper in [16].
@@ 3 Hodge -dual minimal surface
Now, we are going to define the solvable functional $S[C]$ satisfying both the duality and additivity. We are following the general construction from [15], but here we correct an error made in that paper: we only need three components of the quaternionic field $X_{\mu}$, not four111the fourth component was worse than unnecessary, it canceled the correct area derivative, produced by the first three components. In the general proofs in that paper, this cancellation was overlooked. The intended results are restored by dropping the fourth component..
@@ 3.1 The surface
$\delta S=-2\int d\theta\frac{\delta C_{\mu}\Sigma_{\mu\nu}\partial_{\theta}C_{\nu}}{\sqrt{\Sigma^{2}}}$
$\implies\frac{\delta S}{\delta\sigma_{\mu\nu}}=-2\frac{\Sigma_{\mu\nu}}{\sqrt{\Sigma^{2}}}$ (3.7)
Therefore, the Hodge duality of the area element $\Sigma$ with these boundary conditions for generic field $X^{a}_{\mu}(z,\bar{z})$ leads to the same Hodge duality of the area derivative. Note, negative sign for the area derivative of Hodge-dual surface, which was absent in the ordinary minimal area (plus, the difference in the area element $\Sigma$ which is now Hodge-dual).
This self-duality of area derivative makes our Hodge-dual minimal area a zero mode of the loop diffusion operator, leading to the solution of the MM loop equations [12, 14], as we argued in the next section.
@@
In the previous paper [15], we found exact solutions for the special case of a planar bounding loop and proved two inequalities. Unfortunately, four components of $X$ field instead of three were erroneously used in that work. In this paper, we find a complete solution for the Hodge-dual surface, superseding the results of previous work and correcting these errors.
@@ 3.3 The Virasoro Constraint and Uniformization
The existence of such a coordinate system for any surface topology is guaranteed by the Uniformization Theorem [18]. This theorem ensures that we are always free to set $(f^{\prime})^{2}=0$ by a suitable diffeomorphism. In this gauge, the non-linear geometric area density simplifies exactly to the quadratic Lagrangian:
@@
$\frac{\delta S_{\chi}[C]}{\delta\sigma_{\mu\nu}}=-2\frac{(F_{\mu\nu}+\chi\ast F_{\mu\nu})}{\sqrt{2F_{\mu\nu}^{2}}};$ (3.26)
@@
In presence of the Virasoro constraint (3.21), the problem becomes a challenging one, and it was investigated over the last centuries by great mathematicians, starting with Riemann and Hilbert. We give a brief summary of modern state of this theory in the [16], where this theory will be used in the final stage of the analytic solution.
@@ 3.6 Symmetrization of the area by parity
One may then ask: why do we need the Hodge-dual surface at all? If the area derivative of the symmetrized Hodge-dual surface is equivalent (up to normalization) to that of the Nambu-Goto string, why not claim the Nambu-Goto string itself solves the loop equation as a zero mode? Direct calculation in [15] shows this is not possible. The area derivative of the conventional minimal area is well-defined, but it does not solve the loop equation: it produces in this equation a singular factor $\frac{\delta}{\delta^{2}+\epsilon^{2}}$, depending on the ratio of the two infinitesimal cutoff parameters involved in the definition of the loop variations. Depending on this ratio, we obtain an arbitrary number between zero and infinity instead of the required zero in the loop equation.
@@
$\frac{\delta S_{\chi}}{\delta\sigma}$ $=-2(F+\chi\ast F)/\sqrt{2F^{2}};$
$F_{\mu\nu}$ $=2\Im f_{\mu}\bar{f}_{\nu}$
@@ 3.10 Resulting Geometry
This solution represents (up to rotation with $\eta^{a}$ matrices) a general conformal mapping from the parameter domain $\xi$ directly to the physical plane $(X_{1},X_{2})$, with trivial coordinates in the transverse dimensions $X_{3},X_{4}$. The function $w(\xi)$ completely determines the shape of the planar loop. This planar solution matches the one found in [15] for planar loop.
@@
For this planar solution, we can explicitly verify the zero-mode condition. The area derivative for the planar ansatz reduces to the simple algebraic form
$\delta S_{\chi}/\delta\sigma_{\mu\nu}=-2(F_{\mu\nu}+\chi\ast F_{\mu\nu})/\sqrt{2F^{2}}$ (3.51)
which is manifestly self-dual. Therefore, the condition $\frac{\delta S_{\chi}}{\delta\sigma}=\chi\frac{\delta S_{\chi}}{\delta\sigma}$ holds pointwise. By the Bianchi identity (2.5), this implies $\mathcal{L}S[C]\equiv 0$ without reliance on the general loop calculus of Ref [4]. This confirms that the explicit planar solution is indeed an exact zero-mode of the loop operator. This is, however, a rather trivial zero mode, as the normalized tensor $(F_{\mu\nu}+\chi\ast F_{\mu\nu})/\sqrt{2F^{2}}=\eta^{3,\chi}_{\mu\nu}$ is a universal constant tensor, with vanishing functional derivative.
@@ 4.1 Stationarity and Boundary Variation
The existence of the additive minimal solution is guaranteed by the results of [2] and [19], who established that the disconnected solution minimizes the functional for separated or touching boundaries. Since the surface is an extremum in the full functional space, the bulk variation vanishes identically by the principle of stationary action.
@@
While a connected minimal surface (global minimum) may exist, the QCD Loop Equation requires the additive Goldschmidt solution to satisfy the factorization condition $S[C_{1}\!\!\circ C_{2}]=S[C_{1}]+S[C_{2}]$. The topological barrier proven by [6] ensures that this additive solution is a local minimum, which is all we need for the proof of Hodge-duality.
@@
The continuity of the transition between a single loop with an infinitesimal neck and two touching loops (the self-intersecting limit) is mathematically grounded in the Bridge Principle for stable minimal surfaces [22]. This principle establishes that given two stable minimal surfaces bounded by curves $C_{1}$ and $C_{2}$, one can construct a connected minimal surface bounded by a single loop formed by joining $C_{1}$ and $C_{2}$ with a thin “bridge” of width $\epsilon$. As $\epsilon\to 0$, this connected surface converges smoothly (in the varifold sense) to the union of the original surfaces. Consequently, the Goldschmidt solution we utilize—which effectively treats the area as additive, $\mathcal{A}(C_{1}\cup C_{2})=\mathcal{A}(C_{1})+\mathcal{A}(C_{2})$—represents the correct continuous limit of the loop equation dynamics, avoiding the discontinuous jumps associated with the global minimization of the Plateau problem (such as the catenoid collapse). This ensures that the loop derivative remains well-defined across topology changes. Specifically, because the convergence is smooth in the varifold sense, the functional derivatives required for the loop equation (which probe the surface near the loop $C$) are continuous limits of the derivatives of the connected surface, justifying the application of the differential loop calculus to the additive Goldschmidt functional.
We assume (or conjecture, supported by [2, 19, 22]) that the Goldschmidt/local-minimum branch defines a continuous additive functional on immersed loops with transverse self-intersections, and that the loop-space area derivative exists almost everywhere along each branch. Under this assumption, Theorem 2.1 applies.